# Filtered Corpus Training (FiCT) Shows that Language Models can Generalize from Indirect Evidence

### v1

@@ 1 Introduction
This method enables us to ask whether models can form relevant linguistic generalizations from indirect evidence, or whether they require direct evidence (e.g. examples of constructions during training; Warstadt and Bowman, 2022; Mueller and Linzen, 2023). Furthermore, by carefully controlling for the number of parameters we can investigate the inductive biases of two major LM architectures, Transformers and LSTMs, which allows us to give more detailed answers about the recent successes of Transformer models on a fine-grained linguistic level.
@@ 2.1 Surprisal Theory
Language modeling performance can measured using perplexity, indicating a model’s fit to a corpus distribution. Intuitively, one might expect that lower perplexity leads to more human-like linguistic behavior. This connection has been explored in detail in the context of surprisal theory (Hale, 2001; Levy, 2008): encountering a highly surprising token results in a longer reading time. Initial findings indicate that lower perplexity, as measured by language models, leads to better reading time predictions (Fossum and Levy, 2012; Goodkind and Bicknell, 2018; Wilcox et al., 2020), although affected by model architecture (Hao et al., 2020), cross-lingual effects (Kuribayashi et al., 2021), and syntactic ambiguity (Arehalli et al., 2022). It has been shown, however, that lower perplexity only results in better predictive power up to around 2 billion training tokens (Oh and Schuler, 2023a): after this point LMs become too powerful at predicting low-frequency constructions (Oh et al., 2024). The present paper also explores the connection between perplexity and human-like linguistic behavior and will find a dissociation with perplexity.
@@ 2.2 Targeted Syntactic Evaluations
Perplexity should be augmented with other evaluations that specifically target the models’ ability to generalize in a human-like way. Such investigations often draw on psycholinguistic paradigms, treating language models as participants in order to learn what such models “know” about specific linguistic phenomena (Futrell et al., 2019; Ettinger, 2020). A common paradigm in this body of literature, usually referred to as “targeted syntactic evaluations” (Linzen et al., 2016; Jumelet and Hupkes, 2018; Marvin and Linzen, 2018; Gauthier et al., 2020; Newman et al., 2021) involves comparing language models’ preferences between minimal pairs of sentences: a model is deemed to understand a phenomenon if it assigns a higher probability to the grammatical alternation.
The benchmark with the widest coverage over linguistic phenomena is the Benchmark of Linguistic Minimal Pairs (BLiMP, Warstadt et al., 2020). BLiMP consists of 67 different benchmarks, each consisting of 1,000 minimal pairs, which target twelve different linguistic areas, broadly construed, across morphology, syntax, semantics, and the syntax-semantics interface. This is the benchmark we use as a primary means of evaluation in the present investigation, discussed in greater detail in §4.
@@ 2.3 Linguistic Generalization
While targeted syntactic evaluations give an insight into a model’s linguistic competence, it does not show how a model acquires this notion of grammaticality. In this paper we focus on two kinds of linguistic generalizaton. Structural generalization (Hupkes et al., 2023) asks: can language models make grammaticality judgments in syntactically more complex constructions than seen during training? One line of work approaches this question from a fine-tuning perspective: by fine-tuning a model on particular set of constructions we can measure the impact this has on other linguistic constructions (Prasad et al., 2019; Weber et al., 2024). Lexical generalization asks whether models can generalize a seen construction to new lexical items that it has not seen in that construction (Kim and Linzen, 2020).
In order to gain a causal perspective on how the training data influences model performance, we retrain models from scratch on filtered corpora. This methodology has been deployed in earlier work to investigate how LMs learn the licensing conditions of negative polarity items from different contexts (Jumelet et al., 2021; Weber et al., 2021). Warstadt (2022) investigates the poverty of the stimulus debate through the lens of filtered corpora, focusing on the phenomenon of subject auxiliary inversion. Finally, Misra and Mahowald (2024) investigate rare adjective noun constructions and manipulate training corpora to investigate how models acquire an understanding of rare constructions. Whereas most of these focus on a particular linguistic construction, our work applies the approach to a wide range of phenomena.
@@ 3.1 Logic of the Method
In this paper, the FiCT methodology is primarily used to test if LMs are capable of extrapolating linguistic rules learned from environments in training data to unseen environments. In order to ensure that the specified environments are not seen in the training data, we use filters to remove sentences with the specified environments from a naturalistic corpus. By comparing models trained on the ablated data and models trained on the full, naturalistic corpus, we can potentially determine if, how, and when language models are able to make such generalizations.
@@ 3.2.1 Structural Generalization
The benchmark targeted by this filter tests subject-verb number agreement in the presence of an intervening distractor in a prepositional phrase, as illustrated in Figure 1. agr-pp-mod filters all sentences containing the dependency structure $\textsc{verb}>_{\text{nsubj}}\textsc{noun}>_{\text{nmod}}\textsc{noun}>_{\text{case}}\textsc{adp}$. The resulting filtered corpus will still contain PPs modifying nouns in other contexts (e.g. object position). If a learner has formed a general ‘rule’ for subject-verb agreement, and seen PP-modified objects, it should be able to generalize to agreeement with PP-modified subjects, even when it hasn’t seen them during training.
@@
One of the filters targeting determiner-noun agreement focuses on cases where and adjective occurs between a demonstrative determiner and a noun, e.g. These/*This red cars. We create a filter that removes all occurrences of a demonstrative determiner followed by an adjective and a noun. A model can then still infer the number agreement from determiner/noun pairs without an intervening adjective.
@@ 3.2.2 Lexical Generalization
The four BliMP benchmarks targeted by agr-re-irr-sv all test language model performance on subject-verb agreement, targeting “regular” plurals, like dress/dresses and “irregular” plurals, like goose/geese. The filter removes all sentences with nominal subjects where the noun occurs in any of the four benchmarks. A learner on this filtered corpus can still solve the benchmark if it develops a notion of grammatical number, a representation of the grammatical number of the nouns in the benchmark based on their usage in other contexts, and then generalizes the subject-verb agreement it sees for other nouns to these nouns.
@@
Passive constructions can only be formed by transitive verbs. BLiMP targets this phenomenon by replacing transitive verbs in passive constructions by intransitive verbs: John is insulted by Mary vs. *John is smiled by Mary. Much like agr-re-irr-sv and det-noun, the passive filter operates by removing sentences that contain words on a word list in a specific linguistic environment. Concretely, this word list consists of the verbs that are actually used in these two benchmarks in passive form, and the filter removes sentences where such words appear in passive voice.
@@ 4 Experimental Setup
Two architectures are used for the models trained in this investigation: LSTMs (Hochreiter and Schmidhuber, 1997)) and decoder-only Transformers (Vaswani et al., 2017). For each architecture we train separate models on the 16 training corporate for five random seeds each, resulting in a total of 160 models. Model hyperparameters were selected to control for number of parameters as closely as possible. The LSTMs have two layers with embedding and hidden dimension of 1024. Output and embedding layer weights were tied, and we used dropout of 0.1 during training. The Transformers were constructed with feed-forward and hidden layer dimensions of 768, eight attention heads, and eight hidden layers. The LSTMs and the Transformers had 68.0M and 67.1M trainable parameters, respectively.
@@
However, the FICT methodology’s main advantage lies not in looking the performance of each model in isolation, but on the difference in performance between two models that are otherwise identical but for their training data. Thus, for each model and each BLiMP benchmark, a change score (or delta) was calculated with respect to the average performance of all models of the same architecture trained on the full corpus (i.e. average over the five seeds).
@@
$\textsf{acc}\Delta(M,F,B):=\textsf{acc}^{M_{F}}_{B}-\overline{\textsf{acc}^{M_{\textsc{full}}}_{B}}$
@@
Our final evaluation metric looks at the probability deltas between grammatical and ungrammatical sentences, roughly the magnitude of the accuracy metric:
$P\Delta(M,F)(s)=\log P_{M_{F}}(s^{+})-\log P_{M_{F}}(s^{-})$
@@ 5.1 Perplexity
We found that Transformers uniformly achieve lower perplexities on the test corpus than the LSTMs for all training corpora, as expected. The mean test perplexity across all corpora and random seeds was $47.13$ for the Transformers and $53.56$ for the LSTMs; a paired $t$-test of mean perplexities per corpus found the difference between the model types to be significant ($t=270.94$, $p\approx 0$). As noted in §4, while we downsampled all corpora to the same number of lines, the number of tokens varies between different training corpora. Previous research has shown a clear negative relationship between the number of tokens seen in training and test corpus perplexity. This effect is also present in our data, for both architectures (LSTMs: Pearson’s $r=-0.970$; Transformers: $r=-0.976$).
@@ 5.2 TSE Accuracy on BLiMP
Mean overall accuracy on all of BLiMP across different training corpora (i.e. $\overline{\textsf{acc}_{\textsc{all}}^{M_{F}}}$) was $70.4$ for the LSTMs and $71.9$ for the Transformers. This result was statistically significant (paired $t=-17.38$, $p\approx 0$). Figure 5 in Appendix B shows all of the accuracies.
@@ 5.3 Accuracy Delta
A table of the accuracy deltas, averaged across all random seeds, can be found in Figure 2. Mean overall accuracy delta over all benchmarks and across all training corpora (i.e. $\overline{\Delta}(M,F,B)$) was $-0.393$ for the LSTMs and $0.0313$ for the Transformers. This result was statistically significant (paired $t=-5.10$, $p\approx 0.00013$).
Focusing on the $F=F(B)$ cases (i.e. black-outlined cells in the chart), we note that most deltas are generally negative but fairly close to zero, with a few outliers, such as the models trained on the existential-there, agr-pp-mod, and npi-only corpora. These results suggests that, overall, learners are usually able to use various sorts of indirect evidence to acquire correct grammatical generalizations when direct evidence has been made unavailable, as otherwise we could expect much larger deltas across the board.
@@ 5.4 Probability Delta
In order to gain a more fine-grained insight into the impact of corpus filtering, we examine the results at an item-level. For this we make use of the $P\Delta$ metric, which expresses a model’s magnitude of a grammaticality judgment. In Figure 4A we plot the average $P\Delta$ scores for the full models for each BLiMP benchmark, averaged across seeds. It can be seen here that the Transformers and LSTMs result in highly similar $P\Delta$’s ($r=0.98;p\approx 0$), although the Transformer scores are slightly higher on average those of the LSTMs ($2.99$ vs. $2.41$, respectively), which is in line with the significant difference in TSE accuracy of §5.2.
@@
To investigate this observation across all benchmarks we plot the difference in $P\Delta$ going from full to Filtered in Figure 4B. This difference represents the absolute impact of filtering on the TSE task. By plotting the Transformer results against the LSTM we gain an insight whether filtering has a similar impact on both architectures. We observe a strong correlation between these $P\Delta$ differences ($r:0.91$, $p\approx 0$). Subtle difference are present, however, for a number of filters the $P\Delta$ score increases after filtering which is especially prevalant for the Transformer models.
@@ 6 Discussion
Our findings contribute to a growing body of research that suggest a dissociation between perplexity and more targeted evaluations of linguistic competence in artificial learners (Hu et al., 2020). In a carefully controlled setting and for a wide range of phenomena, we demonstrate that the training objective of minimizing perplexity does not predict linguistic generalization. This raises interesting questions on the relation between perplexity and grammaticality judgments (Lau et al., 2017): while Transformers are better at memorizing the structure of its training data, we show they are less capable than LSTMs of forming robust linguistic generalizations. An interesting step for future work would be to uncover what language modeling aspects Transformers do excel at, which allows them to obtain a superior test perplexity.
@@
Where earlier work has focused on specific linguistic constructions, such as subject auxiliary inversion (Warstadt, 2022), relative clauses (Prasad et al., 2019), and negative polarity items (Jumelet et al., 2021; Weber et al., 2021), the results of this paper essentially confirm a similar result for a much wider array of syntactic and semantic phenomena. While in many cases the ablations we performed did clearly negatively affect the performance of our artificial learners on the relevant linguistic evaluations, the magnitude of this effect was generally quite small for all but a small handful of the linguistic phenomena we analyzed. In general, even when tested on the specific benchmarks corresponding to the environments that were ablated from their input, models still perform considerably better than chance. Thus, our research provides evidence in favor of the indirect evidence hypothesis.
@@
As just discussed, the primary contribution of this paper has been the development of the FiCT method and the use of it to demonstrate LMs’ successful generalization from indirect evidence across a wide range of linguisitc phenomena. This success raises a very natural follow-up question: what explains this successful generalization behavior?

### v2

@@ 1 Introduction
This method enables us to ask whether models can form relevant linguistic generalizations from indirect evidence, or whether they require direct evidence (e.g. examples of constructions during training; Warstadt and Bowman, 2022; Mueller and Linzen, 2023). In essence, by intervening on patterns in the training data we obtain a more causal account of the relation between training data and model behavior (Pearl, 2009). Furthermore, by carefully controlling for the number of parameters we can investigate the inductive biases of two major LM architectures, Transformers and LSTMs, which allows us to give more detailed answers about the recent successes of Transformer models on a fine-grained linguistic level.
@@
These results shed light on the debate between memorization and generalization in language models: by causally intervening on the training data, we ensure that models have never seen instances of their evaluation targets. That they can still make correct grammaticality judgments shows they generalize in subtle and linguistically-relevant ways that go beyond their training data.
@@ 2.1 Surprisal Theory
Language modeling performance can be measured using perplexity, indicating a model’s fit to a corpus distribution. Intuitively, one might expect that lower perplexity leads to more human-like linguistic behavior. This connection has been explored in detail in the context of surprisal theory (Hale, 2001; Levy, 2008): encountering a highly surprising token results in a longer reading time. Initial findings indicate that lower perplexity, as measured by language models, leads to better reading time predictions (Fossum and Levy, 2012; Goodkind and Bicknell, 2018; Wilcox et al., 2020), although affected by model architecture (Hao et al., 2020), cross-lingual effects (Kuribayashi et al., 2021), and syntactic ambiguity (Arehalli et al., 2022). It has been shown, however, that lower perplexity only results in better predictive power up to around 2 billion training tokens (Oh and Schuler, 2023a): after this point LMs become too accurate at predicting low-frequency constructions and long-distance dependencies (Oh et al., 2024). The present paper also explores the connection between perplexity and human-like linguistic behavior and will find a dissociation with perplexity.
@@ 2.2 Targeted Syntactic Evaluations
Perplexity should be augmented with other evaluations that specifically target the models’ ability to generalize in a human-like way. Such investigations often draw on psycholinguistic paradigms, treating language models as participants in order to learn what such models “know” about specific linguistic phenomena (Futrell et al., 2019; Warstadt et al., 2019b; Ettinger, 2020). A common paradigm in this body of literature, usually referred to as “targeted syntactic evaluations” (Linzen et al., 2016; Jumelet and Hupkes, 2018; Marvin and Linzen, 2018; Kann et al., 2019; Newman et al., 2021) involves comparing language models’ preferences between minimal pairs of sentences: a model is deemed to understand a phenomenon if it assigns a higher probability to the grammatical alternation.
The benchmark suites with the widest coverage over linguistic phenomena are SyntaxGym (Gauthier et al., 2020) and the Benchmark of Linguistic Minimal Pairs (BLiMP, Warstadt et al., 2020), the latter of which we will use in our experiments. BLiMP consists of 67 different benchmarks, each consisting of 1,000 minimal pairs, which target twelve different linguistic areas, broadly construed, across morphology, syntax, semantics, and the syntax-semantics interface. This is the benchmark we use as a primary means of evaluation in the present investigation, discussed in greater detail in §4.
@@ 2.3 Linguistic Generalization
While targeted syntactic evaluations give an insight into a model’s linguistic competence, it does not show how a model acquires this notion of grammaticality. In this paper we focus on two kinds of linguistic generalization. Structural generalization (Hupkes et al., 2023) asks: can language models make grammaticality judgments in syntactically more complex constructions than seen during training? One line of work approaches this question from a fine-tuning perspective: by fine-tuning a model on a particular set of constructions we can measure the impact that this has on other linguistic constructions (Prasad et al., 2019; Weber et al., 2024). Lexical generalization asks whether models can generalize a seen construction to new lexical items that it has not seen in that construction (Kim and Linzen, 2020).
In order to gain a causal perspective on how the training data influences model performance, we retrain models from scratch on filtered corpora. This methodology has been deployed in earlier work to investigate how LMs learn the licensing conditions of negative polarity items from different contexts (Jumelet et al., 2021; Weber et al., 2021). Warstadt (2022) investigates the poverty of the stimulus debate through the lens of filtered corpora, focusing on the phenomenon of subject auxiliary inversion. Finally, Misra and Mahowald (2024) investigate rare adjective-noun constructions and manipulate training corpora to investigate how models acquire an understanding of rare constructions. Whereas most of these focus on a particular linguistic construction, our work applies the approach to a wide range of phenomena.
@@ 3.1 Logic of the Method
In this paper, the FiCT methodology is primarily used to test whether LMs are capable of extrapolating linguistic rules learned from environments in training data to unseen environments. In order to ensure that the specified environments are not seen in the training data, we use filters to remove sentences with the specified environments from a naturalistic corpus. By comparing models trained on the ablated data and models trained on the full, naturalistic corpus, we can potentially determine whether, how, and when language models are able to make such generalizations.
@@ 3.2.1 Structural Generalization
The benchmark targeted by this filter tests subject-verb number agreement in the presence of an intervening distractor in a prepositional phrase, as illustrated in Figure 1. agr-pp-mod filters all sentences containing the dependency structure $\textsc{verb}>_{\text{nsubj}}\textsc{noun}>_{\text{nmod}}\textsc{noun}>_{\text{case}}\textsc{adp}$. The resulting filtered corpus will still contain PPs modifying nouns in other contexts (e.g. object position). If a learner has formed a general ‘rule’ for subject-verb agreement, and seen PP-modified objects, it should be able to generalize to agreement with PP-modified subjects, even when it hasn’t seen them during training.
@@
One of the filters targeting determiner-noun agreement focuses on cases where an adjective occurs between a demonstrative determiner and a noun, e.g. These/*This red cars. We create a filter that removes all occurrences of a demonstrative determiner followed by an adjective and a noun. A model can then still infer the number agreement from determiner/noun pairs without an intervening adjective.
@@ 3.2.2 Lexical Generalization
The four BliMP benchmarks targeted by agr-re-irr-sv all test language model performance on subject-verb agreement, targeting regular plurals, like dress/dresses and irregular plurals, like goose/geese. The filter removes all sentences with nominal subjects where the noun occurs in any of the four benchmarks. A learner on this filtered corpus can still beat the benchmark if it develops a notion of grammatical number, a representation of the grammatical number of the nouns in the benchmark based on their usage in other contexts, and then generalizes the subject-verb agreement it sees for other nouns to these nouns.
@@
In English, passive constructions can only be formed from transitive verbs. BLiMP targets this phenomenon by replacing transitive verbs in passive constructions by intransitive verbs: John is insulted by Mary vs. *John is smiled by Mary. Much like agr-re-irr-sv and det-noun, the passive filter operates by removing sentences that contain words on a word list in a specific linguistic environment. Concretely, this word list consists of the verbs that are actually used in these two benchmarks in passive form, and the filter removes sentences where such words appear in passive voice.
@@ 4 Experimental Setup
Two architectures are used for the models trained in this investigation: LSTMs (Hochreiter and Schmidhuber, 1997)) and decoder-only Transformers (Vaswani et al., 2017). For each architecture, we train separate models on the 16 training corpora for five random seeds each, resulting in a total of 160 models. Model hyperparameters were selected to control for number of parameters as closely as possible. The LSTMs have two layers with embedding and hidden dimension of 1024. Output and embedding layer weights were tied, and we used dropout of 0.1 during training. The Transformers were constructed with feed-forward and hidden layer dimensions of 768, eight attention heads, and eight hidden layers. The LSTMs and the Transformers had 68.0M and 67.1M trainable parameters, respectively.
@@
However, the FICT methodology’s main advantage lies not in looking at the performance of each model in isolation, but on the difference in performance between two models that are otherwise identical but for their training data. Thus, for each model and each BLiMP benchmark, a change score (or delta) was calculated with respect to the average performance of all models of the same architecture trained on the full corpus (i.e. average over the five seeds).
@@
$\textsf{acc}\Delta(M,F,B):=\textsf{acc}^{M_{F}}_{B}-\overline{\textsf{acc}^{M_{\textsc{full}}}_{B}}$ (1)
@@
Our final evaluation metric looks at the probability deltas between grammatical and ungrammatical sentences:
$P\Delta(M,F)(s)=\log P_{M_{F}}(s^{+})-\log P_{M_{F}}(s^{-})$ (2)
$P\Delta$ expresses the magnitude of a model’s grammaticality judgment: whereas $\textsf{acc}\Delta$ only expresses the ratio of items for which a model assigned a higher probability to the grammatical case, $P\Delta$ can be interpreted as the confidence of a model’s judgment.
@@ 5.1 Perplexity
We found that Transformers uniformly achieve lower perplexities on the test corpus than the LSTMs for all training corpora, as expected. The mean test perplexity across all corpora and random seeds was $47.13$ for the Transformers and $53.56$ for the LSTMs; a paired $t$-test of mean perplexities per corpus found the difference between the model types to be significant ($t=270.94$, $p\ll 0.01$). As noted in §4, while we downsampled all corpora to the same number of lines, the number of tokens varies between different training corpora. Previous research has shown a clear negative relationship between the number of tokens seen in training and test corpus perplexity (Kaplan et al., 2020). This effect is also present in our data, for both architectures (LSTMs: Pearson’s $r=-0.970$; Transformers: $r=-0.976$).
We also investigate the perplexity on the BLiMP sentences for the full and Filtered models. This provides us insight into the likelihood of these sentences: if the model assigns a relatively low likelihood to them, then grammaticality judgments will be less reliable as well (Newman et al., 2021). In Figure 3 we show the scores for this. Surprisingly, the LSTM models yield lower perplexity on the BLiMP sentences than the Transformers. This shows that Transformers have shifted their probability mass to other sentence types than found in BLiMP, but where to exactly remains an open question. Nonetheless, the perplexity scores on BLiMP are similar to the average perplexity on the test corpus, which demonstrates that these items are of similar likelihood.
Figure 3: Perplexity scores on the test corpus ($C^{\textit{test}}$) and the grammatical and ungrammatical BLiMP sentences ($s^{+}$ $s^{-}$). BLiMP scores for the full models are averaged over all benchmarks, and for the Filtered models for their corresponding benchmark.
@@ 5.2 TSE Accuracy on BLiMP
Mean overall accuracy on all of BLiMP across different training corpora (i.e. $\overline{\textsf{acc}_{\textsc{all}}^{M_{F}}}$) was $70.4$ for the LSTMs and $71.9$ for the Transformers. This result was statistically significant (paired $t=-17.38$, $p\ll 0.01$). Figure 6 in Appendix B shows all of the accuracies.
@@ 5.3 Accuracy Delta
A table of the accuracy deltas, averaged across all random seeds, can be found in Figure 2. Mean overall accuracy delta over all benchmarks and across all training corpora (i.e. $\overline{\Delta}(M,F,B)$) was $-0.393$ for the LSTMs and $0.0313$ for the Transformers. This result was statistically significant (paired $t=-5.10$, $p\ll 0.01$).
Focusing on the $F=F(B)$ cases (i.e. black-outlined cells in the chart), we note that most deltas are generally negative but fairly close to zero, with a few outliers, such as the models trained on the existential-there, agr-pp-mod, and npi-only corpora. These results suggest that, overall, learners are usually able to use various sorts of indirect evidence to acquire correct grammatical generalizations when direct evidence has been made unavailable, as otherwise we could expect much larger deltas across the board.
@@ 5.4 Probability Delta
In order to gain a more fine-grained insight into the impact of corpus filtering, we examine the results at an item-level. For this we make use of the $P\Delta$ metric, which expresses a model’s magnitude of a grammaticality judgment. In Figure 5A we plot the average $P\Delta$ scores for the full models for each BLiMP benchmark, averaged across seeds. It can be seen here that the Transformers and LSTMs result in highly similar $P\Delta$’s ($r=0.98;p\approx 0$), although the Transformer scores are slightly higher on average than those of the LSTMs ($2.99$ vs. $2.41$, respectively), which is in line with the significant difference in TSE accuracy of §5.2.
@@
To investigate this observation across all benchmarks we plot the difference in $P\Delta$ going from full to Filtered in Figure 5B. This difference represents the absolute impact of filtering on the TSE task. By plotting the Transformer results against the LSTM we gain an insight whether filtering has a similar impact on both architectures. We observe a strong correlation between these $P\Delta$ differences ($r:0.91$, $p\approx 0$). Subtle difference are present, however, for a number of filters the $P\Delta$ score increases after filtering which is especially prevalent for the Transformer models.
@@ 6 Discussion
Our findings contribute to a growing body of research that suggest a dissociation between perplexity and more targeted evaluations of linguistic competence in artificial learners (Hu et al., 2020). In a carefully controlled setting and for a wide range of phenomena, we demonstrate that the training objective of minimizing perplexity does not predict linguistic generalization. This raises interesting questions on the relation between perplexity and grammaticality judgments (Lau et al., 2017): while Transformers are better at memorizing the structure of its training data, we show they are less capable than LSTMs of forming robust linguistic generalizations. An interesting step for future work would be to uncover what language modeling aspects Transformers do excel at, which allows them to obtain a superior test perplexity (e.g. word frequency, as studied in Wei et al., 2021). Future work should also compare our measure(s) of generalization with others in the literature, given evidence that these are not always well-correlated with each other (Sun et al., 2023).
We also note that while likelihood judgments do not necessarily directly measure grammaticality, since likelihood is the outcome of many other factors (e.g. semantic plausibility, pragmatic felicity), the use of minimal pairs for BLiMP does help control for this since it reports judgments on sentences which differ on (usually) one word, thus keeping these other components constant between the two sentences. That being said, it would be a worthwhile follow-up to conduct probing experiments to more directly model grammaticality judgments, in the style of e.g. Jumelet et al. (2021) (see the next subsection as well). [^2] We thank an anonymous reviewer for encouraging us to think about this distinction.
Our results also have consequences for how we think about language model evaluation more broadly: to the extent that we believe that models should be able to generalize from indirect evidence, we cannot rely on perplexity as the sole measure of LM quality but must measure and test for this ability directly.
@@
Where earlier work has focused on specific linguistic constructions, such as subject auxiliary inversion (Warstadt, 2022), relative clauses (Prasad et al., 2019), and negative polarity items (Warstadt et al., 2019a; Jumelet et al., 2021; Weber et al., 2021), the results of this paper essentially confirm a similar result for a much wider array of syntactic and semantic phenomena. While in many cases the ablations we performed did clearly negatively affect the performance of our artificial learners on the relevant linguistic evaluations, the magnitude of this effect was generally quite small for all but a small handful of the linguistic phenomena we analyzed. In general, even when tested on the specific benchmarks corresponding to the environments that were ablated from their input, models still perform considerably better than chance. Thus, our research provides evidence in favor of the indirect evidence hypothesis.
@@
As just discussed, the primary contribution of this paper has been the development of the FiCT method and the use of it to demonstrate LMs’ successful generalization from indirect evidence across a wide range of linguistic phenomena. This success raises a very natural follow-up question: what explains this successful generalization behavior?